A stuntman jumps off a roof 375ft from the ground. How long will it take him, falling freely, to reach the ground. (use formula s=16t^2)

Answer by stanbon(75887)

A stuntman jumps off a roof 375ft from the ground. How long will it take him, falling freely, to reach the ground. (use formula s=16t^2)
-----
16t^2 = 375
t = sqrt(375/16)
t = (5/4)sqrt(15)
t = 4.84.. seconds
========================
Cheers,
Stan H.